OpenStudy (anonymous):

√(x+1)-(x-1) Could someone explain how that equation simplifies to 2√x?

OpenStudy (calculusfunctions):

Is it\[\sqrt{(x +1)}-(x -1)\]

OpenStudy (anonymous):

I typed the question wrong √(x+1)^2-(x-1)^2 Could someone explain how that equation simplifies to 2√x

OpenStudy (calculusfunctions):

I thought as much!

OpenStudy (calculusfunctions):

So it's\[\sqrt{(x +1)^{2}-(x -1)^{2}}\]Is this correct?

OpenStudy (anonymous):

Yes

OpenStudy (calculusfunctions):

OK so do you know how to simplify the expression inside the root? You have two options: i). expand and collect like terms, or ii). factor by difference of squares method. Choose a method.

OpenStudy (anonymous):

I'm okay with both. Whatever simpler to explain would be nice.

OpenStudy (calculusfunctions):

Both are simple. your choice.

OpenStudy (anonymous):

Collect like terms please then

OpenStudy (calculusfunctions):

OK then, do you know how to do that?

OpenStudy (anonymous):

Is 4x correct?

OpenStudy (calculusfunctions):

Yes! so then we have\[\sqrt{4x}\]Do you know the next step to simplify this?

OpenStudy (anonymous):

Sqrt 2*2*x?

OpenStudy (calculusfunctions):

\[\sqrt[n]{ab}=\sqrt[n]{a}\sqrt[n]{b}\]

OpenStudy (calculusfunctions):

So then by this rule\[\sqrt{4x}=\sqrt{4}\sqrt{x}\]Correct?

OpenStudy (anonymous):

Oh~ I see!

OpenStudy (anonymous):

Thank you so much for your kind help!

OpenStudy (calculusfunctions):

Very Welcome!